OpenStudy (anonymous):

Find the average rate of change of g(x)=−x2+3x+3 from x=2 to x=6 Simplify your answer as much as possible.

OpenStudy (raffle_snaffle):

plug in x = 6, what is g(x) = ?

OpenStudy (anonymous):

g(6) = 57 g(2) =13

OpenStudy (anonymous):

correct?

OpenStudy (phi):

how did you get g(2) = 13 ? can you show what you did ?

OpenStudy (phi):

oh, you ignored the minus sign in front of - x^2 ?

OpenStudy (anonymous):

\[g(x)= -(2)^2 + 3(2) + 3\] \[g(x)= 4+6+3 \] \[g(x)= 10+3 = 13\]

OpenStudy (raffle_snaffle):

Once you find g(2) and g(6) you can find the average rate of change given by Average rate of change = g(6) - g(2)

OpenStudy (phi):

except the minus sign is not inside the square it is -1 * x^2 or -1 * 2*2 = -4

OpenStudy (anonymous):

44?

OpenStudy (phi):

Here is a graph of your problem You have to find the slope of the line